# Multidimensional Rubric-oriented Reward Model Learning via Geometric Projection Reference Constraints

### v1

@@ Abstract
The integration of large language models (LLMs) into medical practice holds transformative potential, yet their real-world clinical utility remains limited by critical alignment challenges: (1) a disconnect between static evaluation benchmarks and dynamic clinical cognitive needs, (2) difficulties in adapting to evolving, multi-source medical standards, and (3) the inability of conventional reward models to capture nuanced, multi-dimensional medical quality criteria. To address these gaps, we propose MR-RML (Multidimensional Rubric-oriented Reward Model Learning) via GPRC (Geometric Projection Reference Constraints), a novel alignment framework that integrates medical standards into a structured "Dimensions-Scenarios-Disciplines" matrix to guide data generation and model optimization. MR-RML introduces three core innovations: (1) a "Dimensions-Scenarios-Disciplines" medical standard system that embeds domain standards into the full training pipeline; (2) an independent multi-dimensional reward model that decomposes evaluation criteria, shifting from real-time rubric-based scoring to internalized reward modeling for improved consistency and cost-efficiency; (3) geometric projection reference constraints that transform medical cognitive logic into mathematical regularization, aligning scoring gradients with clinical reasoning and enabling synthetic data-driven training. Through extensive evaluations on the authoritative medical benchmark Healthbench, our method yields substantial performance gains over the base LLM Qwen-32B (45% on the full subset and 85% on Hard subset, respectively). It achieves a SOTA among open-source LLMs with scores of 62.7 (full subset) and 44.7 (hard subset), while also outperforming the majority of closed-source models.
@@ 1 Introduction
  • Misalignment Between Evaluation and Medical Cognition Static evaluation systems fail to match the dynamic cognitive needs of medical scenarios. Medical cognitive priorities vary significantly, for example, emergency scenarios require prioritizing emergency level judgment and timeliness of risk prompting, while chronic disease management focuses on completeness of follow-up recommendations and guidance for patient adherence. Existing evaluation benchmarks mostly adopt fixed weight assignment models, failing to reflect these differences. Meanwhile, evaluation dimensions are overly simplified: mainstream benchmarks such as MedQA (roy2024beyond) use multiple-choice questions to assess medical fact memorization, but cannot cover medical-critical requirements like compliance, communication quality, and reasoning logic. Although HealthBench (arora2025healthbench) advances scenario-based evaluation, it only provides datasets without evaluation-training closed-loop, limiting its ability to guide model optimization or generalize to specific industry standards.
@@
  • Bottlenecks in Medical Standard Adaption for LLMs Medical standards are characterized by dynamic updates, and multi-source heterogeneity-different regions, scenarios, and departments have exclusive medical guidelines, with ethical norms revised annually. Existing solutions struggle to balance adaptability and cost efficiency. Extending Reinforcement Learning with Verifiable Rewards (RLVR) (su2025crossing) to real-world tasks is challenging, as evaluation relies on nuanced multi-criteria judgments rather than binary correctness. Recent works have extend RLVR beyond strictly verifiable domains by using structured rubrics as reward signals for multi-dimensional/subjective tasks. (huang2025reinforcement) constructs a large-scale rubric library but lacks deep integration with authoritative medical standards, and its rule-based scoring has low reusability. RaR (gunjal2025rubrics) framework implements multi-dimensional evaluation via LLM-generated instance-specific rubrics, but suffers from high real-time scoring costs, task-dependent standards, and over-reliance on LLM-as-Judge (jang2025instajudge; son2024llm) during RL training. A common shortcoming is the absence of an independently trained reward model, leading to inconsistent evaluation, poor generalization, or excessive training costs.
@@
  • Dilemmas in Multi-Dimensional Reward Model Training Traditional Reward Model Learning (RML) (nika2024reward) face three insurmountable limitations in medicine: extremely high annotation costs requiring senior experts, single-scalar output lacking dimension decomposition (failing regulatory and trust requirements), and absence of mathematical constraints (resulting in irrational scoring gradients and weak generalization). Existing loss functions optimize only a few dimensions, unable to adapt to the multi-orthogonal requirements of medical accuracy, safety, and ethical compliance, while ignoring semantic space geometric relationships between samples.
@@
  • RL scoring paradigm innovation: Medical standard knowledge is internalized into a multi-dimensional reward model, enabling refined decomposition and controllable optimization of evaluation dimensions. This shifts the RLVR or real-time rubric-level scoring for RL to "multi-dimensional internalized reward modeling", improving professionalism, consistency, and generalization at lower costs.
@@
This study aims to provide a "medical cognition-oriented, standard-extensible, and cost-controllable" development solution for medical LLMs, promoting medical AI from technically feasible to medically trustworthy. Extensive evaluations on the authoritative medical benchmark Healthbench demonstrate that our method delivers significant improvements against base LLM model and achieve the SOTA performance, as well as surpassing the majority of closed-source models. [^1] https://huggingface.co/mingpinDZJ/Shanzhi-M1

### v2

@@ Abstract
The integration of large language models (LLMs) into medical practice offers transformative potential, yet their real-world clinical applicability remains constrained by critical alignment issues: (1) a misalignment between static evaluation benchmarks and the dynamic cognitive demands of clinical practice, (2) challenges in adapting to continuously evolving, multi-source medical standards, and (3) the limited capacity of conventional reward models to reflect nuanced, multi-dimensional medical quality criteria. To overcome these limitations, we introduce MR-RML (Multidimensional Rubric-oriented Reward Model Learning) with GPRC (Geometric Projection Reference Constraints)-a novel alignment framework that structured medical standards into a multi-perspective matrix to guide both data generation and model optimization. Our approach introduces three key innovations: (1) a medical standard system that embeds domain-specific guidelines throughout the training pipeline; (2) an independent multi-dimensional reward model that decomposes evaluation criteria, transitioning from rule-based or LLM-based scoring to internalized reward modeling for better evaluation performance; and (3) geometric projection reference constraints that translate clinical cognitive logic into mathematical regularization, aligning scoring gradients with clinical reasoning and facilitating training with synthetically generated data. Extensive evaluations on the authoritative medical benchmark Healthbench demonstrate that our method significantly boosts the performance of the base Qwen-32B model, with improvements of 45% on the full subset and 85% on the hard subset. It achieves state-of-the-art results among open-source LLMs, scoring 62.7 (full) and 44.7 (hard), while also surpassing the majority of closed-source models.
@@ 1 Introduction
  • Misalignment Between Evaluation and Medical Cognition. Static evaluation systems fail to match the dynamic cognitive needs of medical scenarios. Medical cognitive priorities vary significantly, for example, emergency scenarios require prioritizing emergency level judgment and timeliness of risk prompting, while chronic disease management focuses on completeness of follow-up recommendations and guidance for patient adherence. Existing evaluation benchmarks mostly adopt fixed weight assignment models, failing to reflect these differences. Meanwhile, evaluation dimensions are overly simplified: mainstream benchmarks such as MedQA (roy2024beyond) use multiple-choice questions to assess medical fact memorization, but cannot cover medical-critical requirements like compliance, communication quality, and reasoning logic. Although HealthBench (arora2025healthbench) advances scenario-based evaluation, it only provides datasets without evaluation-training closed-loop, limiting its ability to guide model optimization or generalize to specific industry standards.
@@
  • Bottlenecks in Medical Standard Adaption for LLMs. Medical standards are characterized by dynamic updates, and multi-source heterogeneity-different regions, scenarios, and departments have exclusive medical guidelines, with ethical norms revised annually. Existing solutions struggle to balance adaptability and cost efficiency. Extending Reinforcement Learning with Verifiable Rewards (RLVR) (su2025crossing) to real-world tasks is challenging, as evaluation relies on nuanced multi-criteria judgments rather than binary correctness. Recent works have extend RLVR beyond strictly verifiable domains by using structured rubrics as reward signals for multi-dimensional/subjective tasks. (huang2025reinforcement) constructs a large-scale rubric library but lacks deep integration with authoritative medical standards, and its rule-based scoring has low reusability. RaR (gunjal2025rubrics) framework implements multi-dimensional evaluation via LLM-generated instance-specific rubrics, but suffers from high real-time scoring costs, task-dependent standards, and over-reliance on LLM-as-Judge (jang2025instajudge; son2024llm) during Reinforcement Learning (RL). A common shortcoming is the absence of an independently trained reward model, leading to inconsistent evaluation, poor generalization, or excessive training costs.
@@
  • Dilemmas in Multi-Dimensional Reward Model Training. Traditional Reward Model Learning (RML) (nika2024reward) face three insurmountable limitations in medicine: extremely high annotation costs requiring senior experts, single-scalar output lacking dimension decomposition (failing regulatory and trust requirements), and absence of mathematical constraints (resulting in irrational scoring gradients and weak generalization). Existing loss functions optimize only a few dimensions, unable to adapt to the multi-orthogonal requirements of medical accuracy, safety, and ethical compliance, while ignoring semantic space geometric relationships between samples.
@@
  • Reinforcement learning scoring paradigm innovation: Medical standard knowledge is internalized into a multi-dimensional reward model, enabling refined decomposition and controllable optimization of evaluation dimensions. This shifts the RLVR or real-time rubric-level scoring for RL to "multi-dimensional internalized reward modeling", improving professionalism, consistency, and generalization at lower labor-costs.
@@
This study aims to provide a "medical cognition-oriented, standard-extensible, and cost-controllable" development solution for medical LLMs, promoting medical AI from technically feasible to medically trustworthy. Extensive evaluations on the authoritative medical benchmark Healthbench demonstrate that our method delivers significant improvements against base LLM model and achieve the SOTA performance, as well as surpassing the majority of closed-source models. [^1] https://www.modelscope.cn/models/dazhuanjia/Shanzhi-M1